Hank and Sally went shopping for school supplies at the local discount store. Hank realized that the price of two pencils was the same as a blue pen. When checking out, Sally paid $15 for the ten pencils and five pens for Hank and her. What was the price for one pen and the price for one pencil?
My Work so far:
2p = 1b (p=pencil; b=blue pen)
15= 10p + 5b
: Hank and Sally went shopping for school supplies at the local discount store. Hank realized that the price of two pencils was the same as a blue pen. When checking out, Sally paid $15 for the ten pencils and five pens for Hank and her. What was the price for one pen and the price for one pencil?
My Work so far:
2p = 1b (p=pencil; b=blue pen)
15= 10p + 5b

Answer by vleith(1156)

so far so good.
2p = 1b (p=pencil; b=blue pen)
15= 10p + 5b
Now substitute using 2p=b
15 = 10p + 5(2p)
15 = 10P + 10p
15 = 20p
3/4 = p
pencils are $0.75 each
pens are 2p = 2*(0.75) = $1.50
Check your answer.
Is 10*0.75 + 5*(1.50) = 15?